If $4000 is borrowed at a rate of 4.75% interest per year, compounded quarterly, find the amount due at the end of the given number of years. (Round your answers to the nearest cent.)
a) 5 years
b) 7 years
c) 9 years

Answer by lwsshak3(11628)

If $4000 is borrowed at a rate of 4.75% interest per year, compounded quarterly, find the amount due at the end of the given number of years. (Round your answers to the nearest cent.)
a) 5 years
b) 7 years
c) 9 years
***
Compound interest formula: A=P(1+i)^n, P=initial investment, interest rate per period, A=amount after n periods.
For given problem:
P=$4000
i=.0475/4 (quarterly compounding)
n=(4 periods/year), 20, 28, or 36
...
n=20(5 yrs)
A=4000(1+(.0475/4))^20=4000*1.2663=5065.21
...
n=28(7 yrs)
A=4000(1+(.0475/4))^28=4000*1.3917=5566.88
...
n=36(9 yrs)
A=4000(1+(.0475/4))^36=4000*1.5296=6118.25
...
amount due at the end of:
a) 5 years=$5065.21
b) 7 years=$5566.88
c) 9 years=$6118.25